# A Training-Free Style-aligned Image Generation with Scale-wise Autoregressive Model

### v1

@@ 6.1 Additional analysis on generation process
6.2 Limitation of our method
@@
6.3 Analysis on mid-stage pivotal feature interpolation
We conduct additional experiments to assess the impact of applying pivotal feature interpolation at different mid-stage steps (e.g., 3rd to 7th). As illustrated in Fig. 12, the results support our hypothesis that the next-scale prediction process plays a crucial role in determining object placement and overall style. Moreover, the results show that pivotal feature interpolation can be applied at any mid-stage step with minimal performance variation, making it a flexible hyperparameter. This demonstrates the robustness of our model in maintaining style consistency and object alignment across various interpolation points.
7 Additional style-aligned image generation results
@@
Figure 14: Style-aligned image generation results using our model. Each row presents a set of objects rendered in a specific logo style, including tech startup logos, energy logos, digital network logos, cybersecurity logos, and in arid landscape styles.

### v2

@@ B.1 Additional analysis on generation process
B.2 Analysis on mid-stage pivotal feature interpolation
We conduct additional experiments to assess the impact of applying pivotal feature interpolation at different mid-stage steps (e.g., 3rd to 7th). As illustrated in Fig. 11, the results support our hypothesis that the next-scale prediction process plays a crucial role in determining object placement and overall style. Moreover, the results show that pivotal feature interpolation can be applied at any mid-stage step with minimal variation in performance, making it a flexible hyperparameter. This demonstrates the robustness of our model in maintaining style consistency and object alignment across various interpolation points.
Method $S_{\text{dual}}$ $\uparrow$ $S_{\text{obj}}$ $\uparrow$ $S_{\text{sty}}$ $\uparrow$
rr=1 0.364 0.229 0.891
rr=2 0.372 0.247 0.756
rr=3 0.373 0.276 0.577
rr=3.4 0.375 0.282 0.556
rr=4 0.366 0.287 0.504
rr=5 0.356 0.291 0.458
$\alpha_{\text{pivot}}$=0.2 0.370 0.284 0.532
$\alpha_{\text{pivot}}$=0.3 0.371 0.283 0.536
$\alpha_{\text{pivot}}$=0.4 0.375 0.282 0.556
$\alpha_{\text{pivot}}$=0.5 0.372 0.277 0.567
$\alpha_{\text{pivot}}$=0.6 0.373 0.276 0.601
Table 6: Additional ablation study on hyperparameter decay rate $r$ and pivot weight $\alpha_{\text{pivot}}$. Ours.
B.3 Hyperparameter robustness
We also performed ablations on the decay rate $r$ and the pivot weight $\alpha_{\text{pivot}}$ in Tab. 6. The results demonstrate that our method is robust to variations in both $r$ and $\alpha_{\text{pivot}}$ in terms of dual consistency $S_{\text{dual}}$. Additionally, we observe a trade-off between object relevancy and style consistency, highlighting the importance of balancing these aspects. Thus, we highlight that our method offers flexible control—appropriate hyperparameter tuning enables high object relevancy while preserving strong dual consistency.
B.4 Quantitative analysis of our observation
To further support our design choices in Sec. 3.1 of the main paper, we present a quantitative analysis of the scale-wise generation process, focusing on feature dynamics across early, mid, and late stages.
Specifically, we measure the average differences between consecutive generation steps in both content and style features. The content features are extracted using VGG19 activations, while the style features are obtained via DINO ViT-B/8 embeddings. For each region of the generation process—early (steps 1–2), mid (steps 3–7), and late (steps 8–12)—we compute the mean feature changes as follows:
  • Content feature differences: (Early: 0.031, Mid: 0.137, Late: 0.042)
  • Style feature differences: (Early: 0.002, Mid: 0.158, Late: 0.018)
These statistics indicate that both content and style features undergo significantly greater variation during the mid-stage compared to the early or late stages. This validates our hypothesis that the mid-stage plays a pivotal role in defining the spatial layout and visual style of the generated images.
Building on this observation, we propose pivotal feature interpolation and dynamic style injection specifically at the mid-stage to reinforce style alignment and ensure coherent object placement across the generated image set. The effectiveness of this strategy is further confirmed by both quantitative and qualitative results presented in the main paper and this appendix.
B.5 Limitation of our method
@@
Metric Ours FireFlow RF-Solver-Edit RF-Inversion
$S_{\text{obj}}$ $\uparrow$ 0.282 0.281 0.302 0.267
$S_{\text{sty}}$ $\uparrow$ 0.556 0.530 0.292 0.425
$S_{\text{dual}}$ $\uparrow$ 0.375 0.367 0.270 0.328
Time (s) $\downarrow$ 1.58 11.25 633.20 544.06
Table 7: Comparison with recent image editing models.
Appendix C Comparison to recent image editing models
We further conducted comparisons with recent image editing models, including FireFlow [deng2024fireflow], RF-Solver-Edit [wang2024taming], and RF-Inversion [rout2024semantic], to demonstrate the effectiveness of our model. As shown in Tab. 7, our model consistently outperforms existing methods. Moreover, unlike exemplar-based methods that require reference images, our text-driven approach achieves strong performance while offering greater flexibility and practicality.
Appendix D Additional style-aligned image generation results
@@
Figure 14: Style-aligned image generation results using our model. Each row presents a set of objects rendered in a specific logo style, including tech startup logos, energy logos, digital network logos, cybersecurity logos, and music festival logos styles.